At a club play, 117 ticets were sold. Adults' tickets cost $1.25, and children's tickets cost $0.75. In all, $129.75 was taken in. How many of each kind of ticket was sold?
Answer by checkley77(12569)

A+C=117 OR A=117-C
1.25A+.75C=129.75
1.25(117-C)+.75C=129.75
146.25-1.25C+.75C=129.75
-.5C=129.75-146.25
-.5C=-16.50
C=-16.50/-.5
C=33 CHILDREN TICKETS SOLD.
117-33=84 ADULT TICKETS SOLD.
PROOF:
1.25*84+.75*33=129.75
105+24.75=129.75
129.75=129.75